### v1

# HEIE: MLLM-Based Hierarchical Explainable AIGC Image Implausibility Evaluator

Abstract
AIGC images are prevalent across various fields, yet they frequently suffer from quality issues like artifacts and unnatural textures. Specialized models aim to predict defect region heatmaps but face two primary challenges: (1) lack of explainability, failing to provide reasons and analyses for subtle defects, and (2) inability to leverage common sense and logical reasoning, leading to poor generalization. Multimodal large language models (MLLMs) promise better comprehension and reasoning but face their own challenges: (1) difficulty in fine-grained defect localization due to the limitations in capturing tiny details; and (2) constraints in providing pixel-wise outputs necessary for precise heatmap generation. To address these challenges, we propose HEIE: a novel MLLM-Based Hierarchical Explainable image Implausibility Evaluator. We introduce the CoT-Driven Explainable Trinity Evaluator, which integrates heatmaps, scores, and explanation outputs, using CoT to decompose complex tasks into subtasks of increasing difficulty and enhance interpretability. Our Adaptive Hierarchical Implausibility Mapper synergizes low-level image features with high-level mapper tokens from LLMs, enabling precise local-to-global hierarchical heatmap predictions through an uncertainty-based adaptive token approach. Moreover, we propose a new dataset: Expl-AIGI-Eval, designed to facilitate interpretable implausibility evaluation of AIGC images. Our method demonstrates state-of-the-art performance through extensive experiments.
Figure 1: (a) Specialized models lack the ability to explain and analyze subtle implausibility regions, hindering understanding for general users. (b) MLLMs struggle with precise localization of local defects and cannot directly output pixel-level implausibility areas. (c) Our CoT-Driven Explainable Trinity Evaluator can generate heatmaps, analyses, and scores. In our Adaptive Hierarchical Implausibility Mapper, local and global heatmaps are predicted separately, improving the localization of tiny implausibilities.
1 Introduction
Recently, Artificial Intelligence Generated Content (AIGC) has rapidly advanced, leading to the widespread use of AIGC images across various fields. However, they often face quality issues such as artifacts, unnatural textures, and problematic regions. Accordingly, many studies focus on the evaluation of AIGC images [44, 42, 19, 26]. Current evaluations primarily yield scalar scores, simple 0-1 values, failing to identify specific defect regions, which hinders targeted improvements in generative models.
A few months ago, RichHF [23] introduced the task of predicting defect region heatmaps and the RAHF model, which locates specific problematic areas. However, specialized models like RAHF have several limitations in the implausibility evaluation task:
  1. 1.
    Lack of explainability. As shown in Fig. 1 (a), implausibility regions in AIGC images can be complex, such as twisted fingers or overly large eye corners. Specialized small models fail to provide detailed explanations for these subtle defects, making it difficult for general users to understand and optimize the AIGC models.
  2. 2.
    Lack of general knowledge. These small specialized models are trained on limited datasets and lack comprehensive world knowledge, common sense, and logical reasoning skills. As a result, they struggle to generalize to out-of-domain samples during practical usage.
One approach to tackle the above issues is to use MLLMs [3, 24, 11]. MLLMs have rapidly advanced through large training data and excel in understanding, generating, and possessing extensive knowledge and common. However, directly applying vanilla MLLM to image implausibility evaluation remains challenging:
  1. 1.
    Fine-Grained Defect Localization: While MLLMs excel in capturing global image structure, they struggle with pinpointing fine-grained defects. Tiny defect areas like the corners of eyes or eyebrows (Fig. 1 (b)) are often neglected by strong MLLMs such as GPT-4o.
  2. 2.
    Pixel-Wise Output Limitations: General-purpose MLLMs handle various modalities (text, images, videos, etc.) but usually output text only, lacking the capability to produce pixel-level defect heatmaps (Fig. 1 (b)).
To overcome the issues of specialized small models and MLLMs in image implausibility evaluation, we propose HEIE: an MLLM-based Hierarchical Explainable image Implausibility Evaluator, which overcomes various problems as follows:
Firstly, to address the issues of insufficient interpretability, common sense, and logical reasoning in previous methods, we propose the CoT-Driven Explainable Trinity Evaluator. By decomposing the complex task into a sequential task structure, progressing from simple to difficult, our Chain of Thought (CoT) framework harnesses the full potential of LLMs to conduct step-by-step reasoning for intricate implausibility assessments. Our trinity evaluator not only generates scores and heatmaps but also incorporates the analysis and explanation of the implausibility. Crucially, through the integration of CoT [36, 20, 51] and structural design, our heatmaps, scores, and explanations are synergistic and mutually reinforcing.
Secondly, to overcome limitations in precise implausibility localization and pixel-level heatmap generation, we introduce the Adaptive Hierarchical Implausibility Mapper. Our Implausibility Mapper, based on two-way cross-attention, combines low-level image features from ViT  [1] with high-level special map tokens from LLMs to predict implausibility heatmaps. AIGC image defects manifest as either global anomalies, such as extra limbs on an animal, or local irregularities, like issues with individual fingers. So, we propose an uncertainty-driven adaptive algorithm that integrates both global and local tokens. Furthermore, the adaptive map token dynamically adjusts according to image resolution and aspect ratio, thereby ensuring both flexibility and optimal performance.
Lastly, current image implausibility datasets do not meet our explainable evaluation needs. We propose a new dataset: Expl-AIGI-Eval (Explainable AIGC Image Implausibility Evaluation Dataset). Utilizing visual prompting, LLM free-form outputting, and in-context learning-based formatting, this dataset provides interpretable implausibility evaluations, including image descriptions, problematic region identifications, and issue analyses, aiding CoT-based reasoning for LLMs.
Our contributions are as follows:
  1. 1.
    We are the first to achieve explainable image implausibility evaluation. Our CoT-Driven Explainable System guides the LLM to decompose complex problems into subproblems of increasing difficulty and mutually reinforcing, improving heatmaps, scores, and explanations.
  2. 2.
    We are the first to use MLLM for image implausibility heatmap prediction, leveraging MLLM’s commonsense, logical reasoning, and generalization capabilities.
  3. 3.
    We propose the Adaptive Hierarchical Implausibility Mapper, employing dynamic special map tokens for uncertainty-based global-local heatmap predictions, enhancing the localization of subtle defects.
  4. 4.
    We construct a high-quality explainable AIGC image implausibility evaluation dataset, which will be publicly available.
Our method achieves state-of-the-art performance on various datasets and tasks. Extensive comparisons and ablation studies validate the effectiveness of our approach.
Figure 2: The Implausibility Mapper processes a dynamic number of special [MAP] tokens and the image features to enhance detailed implausibility localization. We implement the Adaptive Hierarchical Implausibility Mapper through the local and global heatmaps and the uncertainty-based adaptive fusion. In the verisimilitude scorer, features from the heatmap and the special [SCORE] token are integrated for score prediction. Furthermore, our Cot-Driven Explainable System guides the LLM to decompose complex issues into progressive subproblems, facilitating the mutual enhancement of heatmap, analysis, and score prediction, thus improving explainability.
2 Related Work
2.1 AI-Generated Images
Advancements in text-to-image generation models have significantly impacted creative industries like art, design, and advertising. Stable Diffusion [30]uses denoising processes to produce high-quality images. It is open-source, encouraging academic and industrial applications. DALL·E 3 [4] refines the generation of high-resolution, realistic images with detailed text integration. MidJourney [27], by an independent lab, focuses on creating artistic and imaginative images, prioritizing aesthetics over realism. Recent works [22] explore combining diffusion models with GANs, showing potential for enhancing image quality and diversity. However, AIGC images often contain many implausible elements that require careful evaluation.
2.2 AIGC Image Evaluation
Effective evaluation are required to optimize generative models. Many studies have released datasets to emulate human assessment [44, 42, 19, 14, 21]. Large-scale visual understanding models are now commonly used for reliable evaluation due to their strong semantic capabilities [28, 41, 50]. Prior research has shown that MLLMs are proficient in both low-level visual evaluations [37, 39] and high-level semantic comprehension [52, 49]. Some studies  [38, 34, 17, 40] activate the visual evaluation capabilities of multimodal MLLMs through data fine-tuning.
Existing models primarily aim to align their scores with human preferences [38, 42] or aesthetic evaluation  [16, 15]. These approaches fail to identify specific fine-grained problem areas within images, which hinders users and researchers from making targeted improvements.
Recently, HumanRefiner [10] and RichHF  [23] have introduced tasks focused on detecting fine-grained problematic regions in images. HumanRefiner targets defects in AIGC-generated human images, while RichHF encompasses defects in various objects. However, these approaches rely on proprietary small models to detect defect regions. These models are limited in common sense and knowledge, resulting in insufficient generalization capabilities in complex scenarios. Additionally, proprietary small models lack interpretability and cannot provide detailed reasons or explanations  [2, 9] for the detected issues.
3 Method
3.1 Overview
As shown in Fig. 2, HEIE first employs an Adaptive Hierarchical Implausibility Mapper to generate an image implausibility region heatmap. Subsequently, a verisimilitude scorer is used to predict the image score. Furthermore, with the aid of a Co-driven Explainable System, we stimulated the potential of the LLM to achieve mutual enhancement and joint output of the heatmap, score, and explanation.
3.2 Adaptive Hierarchical Implausibility Mapper
3.2.1 MLLM-Based Implausibility Mapper
To obtain fine-grained, precise image implausibility regions, pixel-level mask outputs are necessary. However, most MLLMs only provide text information and cannot produce heatmap predictions. To address this, we design the Image Implausibility Mapper module for heatmap prediction. We introduced a special map token [MAP] to extract implausibility region information from LLM. And we use image features $F$ from the image encoder and [MAP] token features $T$ from the LLM for a two-way cross-attention calculation [18, 1, 35], as shown in Eq. LABEL:2waytrans. To fuse the [MAP] token embeddings with image features, each input alternates as the query and value during cross-attention operations. Specifically, The special map token embeddings $T$ and image features $F$ server as $E_{\text{text}}$ and $E_{\text{img}}$ respectively in the following two-way cross-attention:
$C_{\text{text\textrightarrow img}}=\text{CrossAtten}(\text{SelfAtten}(E_{\text{text}}),E_{\text{img}},E_{\text{img}})$ (1)
$E_{\text{text}^{{}^{\prime}}}=\text{mlp}(\text{norm}(C_{\text{text\textrightarrow img}}))$
$C_{\text{img\textrightarrow text}}=\text{CrossAtten}(E_{\text{img}},E_{\text{text}^{{}^{\prime}}},E_{\text{text}^{{}^{\prime}}})$
$O_{2way\_CA}=\text{norm}(C_{\text{img\textrightarrow text}})$
By stacking the above two-way cross-attention twice, we get the predicted heatmap $H$.
Data imbalance poses a challenge in heatmap prediction, with mean squared error (MSE) loss often leading to conservative and low-value predictions. And we simply adopt focal loss to overcome the data imbalance issue [31].
3.2.2 Hierarchical Implausibility Mapper
Implausibility detection is not the primary objective of mainstream general-purpose MLLMs. Defect localization often requires detailed information, including small defects like fingers and eyes, making it crucial to detect these nuances. We developed a method targeting both global and local defects by segmenting images into adaptive patches according to resolution  [7]. The image encoder processes the thumbnail and $N$ segmented patches, generating global features $F_{g}$ and local features $F_{i}$. The LLM then outputs $N$ local information-enhanced tokens, $\mathbf{[MAP]_{i}}$, for each patch. Tokens $T_{i}$ combine with local features $F_{i}$ and input into the Implausibility Mapper (see Sec. 3.2.1), where cross-attention mechanisms generate local implausibility heatmaps $L_{i}$. These heatmaps are concatenated to form a complete heatmap:
$H_{l}=F_{concat}(L_{1},L_{2},...,L_{N})$ (2)
To handle large implausibility regions and preserve semantics, we predict a global map token $T_{g}$. We utilize global features $F_{g}$ and the LLM token $T_{g}$ as inputs to the Implausibility Mapper to generate the heatmap $H_{g}$.
3.2.3 Adaptive Implausibility Mapper
AIGC images vary in resolution, aspect ratio, and content, necessitating adaptive strategies for handling implausibility.
Adaptive special map token: We propose an adaptive special map token to manage local information for AIGC images. Our hierarchical design adjusts the local token count based on image scale: more tokens for higher resolution, fewer for lower.
Adaptive Heatmap Fuser: Images can have both global and local defects. Global defects, like a person with three legs, require global features for implausibility prediction, while local errors, like incorrect hand or eye details, need local features.
We introduce an uncertainty-based local-global fusion strategy. We model Local Heatmap $H_{l}$ and Global Heatmap $H_{g}$ as Laplace distributions [^1] The Laplace distribution is $f_{X}(x)=\frac{1}{2\lambda}\exp\left(-\frac{|x-\mu|}{\lambda}\right)$, where $\mu$ and $\lambda$ are parameters. The standard deviation $\sigma$ is $\sigma=\sqrt{2}\lambda$. [25], with standard deviations $\sigma_{l}$ and $\sigma_{g}$. During training, we minimize the terms $H$ and $\sigma$, where $H$ represents either $H_{l}$ or $H_{g}$, and $\sigma$ represents $\sigma_{l}$ or $\sigma_{g}$, respectively:
$H^{*},\sigma^{*}=\arg\min_{H,\sigma}\left(\frac{\sqrt{2}}{\sigma}|H-H^{gt}|+\log(\sigma)\right)$ (3)
During inference, we calculate confidence with Laplace-derived uncertainty and perform weighted summation using $p_{uncertainty}=e^{-\sigma}$ to obtain the final heatmap.
Figure 3: Three stages in Expl-AIGI-Eval dataset construction: Stage 1 - Visual Prompting: Defect regions are circled on images to aid Claude-3.5-sonnet in accurately locating problem areas. Stage 2 - LLM Free-Form Output: Claude-3.5-sonnet generates free-form defect location and analysis. Stage 3 - In-Context Learning-Based Formatting: GPT-4o is used for format standardization.
3.3 CoT-Driven Explainable Trinity Evaluator
3.3.1 Verisimilitude Scorer
Our model generates a heatmap of implausible regions and an overall verisimilitude score for comprehensive evaluation. LLMs struggle with numerical outputs, so we propose the Verisimilitude Scorer method to regress the score.
We define a special score token [SCORE] within the LLM, capturing the image’s verisimilitude. The hidden state of the [SCORE] token is processed by a Feed-Forward Network (FFN) to regress an initial score, $S_{token}$.
Our experiments (Tab. 6) show a strong correlation between the heatmap and verisimilitude score. So we employ several convolution and FFN layers to get the heatmap score $S_{map}$ from the predicted heatmap. The final score $S$ is derived by calibrating $S_{token}$ and $S_{map}$: $S=Calib(S_{token},S_{map})$. In our implementation, we explored various calibration functions, such as weighted summation and dynamic fusion. They exhibited similar performance.
3.3.2 CoT-Driven Explainable System
In addition to generating the implausibility region heatmap and verisimilitude score, it is essential to analyze and explain these outputs for AIGC system improvement. We introduce the CoT-Driven Explainable Trinity Evaluator, depicted in Fig. 2. The LLM is guided through a Chain of Thought (CoT) prompting process in five steps:
1. Image Description: The LLM describes the image, capturing key elements in natural language.
2. Problematic Regions Identification: Based on the description and the understanding of the image, the LLM identifies potential problematic areas for further analysis.
3. Special map token: Based on the preceding Problematic Regions Identification, the LLM injects defect region information into the special map token for use by the implausibility mapper.
4. Analysis of Issues: Based on the preceding implausibility localization, the LLM provides detailed text explanations for each region, including type, causes, and so on.
5. Special Score Token: Based on the comprehensive understanding, the LLM incorporates overall information into a special score token for use by our Verisimilitude Scorer.
This CoT-driven process breaks down the complex evaluation task into five progressively challenging and interrelated tasks, fully leveraging the LLM’s potential.
Trinity Implausibility Evaluator. Textual explanations, heatmaps, and scores synergize in image assessment. Texts describe image content, offering semantic context and guiding heatmap generation to highlight defect-prone areas. Heatmaps, using visual saliency, pinpoint critical regions, aiding scores in accurately quantifying defect levels. Conversely, scores refine heatmap focus, enhancing assessment precision. This interaction enables models to better understand and evaluate image issues.
Figure 4: Model outputs of HEIE. HEIE not only predicts implausibility heatmaps but also provides image descriptions, problematic regions identification, analysis of issues, and score, achieving reliable and explainable implausibility evaluation. Note that for the last AIGC image, which has no evident defects, our model avoids false positives.
Figure 5: Comparison with Baselines. Each set of images, from left to right, includes: (a) Input Image, (b) Output of InternViT-300M-448px, (c) Output of CLIP-ViT-Base-Patch16, (d) Output of ours HEIE, (e) Ground Truth.
3.4 Expl-AIGI-Eval Dataset
Existing AIGC image implausibility datasets, such as RichHF-18K [23] and AbHuman [10], include AIGC images and annotated defect regions but lack interpretability explanations. To address this, we introduce Expl-AIGI-Eval: Explainable AIGC Image Implausibility Evaluation. We provide detailed analyses and explanations for each implausibility region. The annotation process is conducted in three stages, shown in Fig. 3.
1. Visual Prompting: Research [6] indicates that LLMs effectively adapt to visual prompts. Consequently, we annotate defect regions directly on images, ensuring visibility while preserving the image’s integrity and facilitating subsequent use by Claude-3.5-sonnet.
2. LLM Free-Form Outputting: Using Claude-3.5-sonnet, we generate detailed analyses from visual prompts. Prior studies [33] reveal that fixed-format enforcement degrades performance, so we allow free-form outputs for defect causes and analyses.
3. In-Context Learning-Based Formatting: To ensure consistency, we employ in-context learning to guide the GPT-4o in refining and formatting information.
This hierarchical approach ensures comprehensive annotations. Each explanation includes image description, problematic regions identification, and issue analysis, supporting our CoT process in Sec. 3.3.2 and aiding CoT training.
In summary, our Expl-AIGI-Eval dataset comprises 42,760 AIGC images divided into two parts: RichHF-18K [23] and AbHuman [10]. We create 20 diverse query expressions per image for LLM instructions. We manually review 20% of the generated content for quality assurance. Sensitive images related to pornography, privacy, animal protection, and political topics are filtered out to avoid ethical issues.
Figure 6: Results of Hierarchical Implausibility Mapper. Columns (c) and (d) are the global and local heatmaps from our Hierarchical Implausibility Mapper. Column (e) illustrates the final heatmap after adaptively fusing the global and local heatmaps.
Heatmap Prediction Plausibility Score
MSE (All Data)↓ MSE (GT=0)↓ KLD↓ CC↑ SIM↑ AUC-Judd↑ PLCC↑ SRCC↑
PickScore (off-the-shelf) - - - - - - 0.010 0.028
EVA-CLIP encoder (fine-tuned) 0.01614 0.00512 2.835 0.350 0.082 0.549 0.157 0.143
CLIP encoder (fine-tuned) 0.01437 0.00425 2.462 0.251 0.122 0.747 0.390 0.378
RAHF (multi-head) 0.01216 0.00141 1.971 0.425 0.302 0.877 0.666 0.654
RAHF (augmented prompt) 0.00920 0.00095 1.652 0.556 0.409 0.913 0.693 0.681
HEIE (ours) 0.00825 0.00014 1.634 0.574 0.417 0.915 0.697 0.683
Table 1: Comparison with State-of-the-Art on the RichHF-18K Dataset.
Perplexity ↓ GPT-4o Eval ↑ Human Eval ↑
Qwen2-VL-7B-Instruct [3] 1.924209 1.910995 1.979058
DeepSeek-VL-7B-chat [24] 1.794179 1.952880 1.883770
InternVL2-8B [8] 1.456884 2.695288 2.603141
GLM-4V-9B [11] 1.320043 2.486911 2.653403
GPT-4o - 3.828272 3.998953
Claude-3.5-Sonnet - 3.938220 4.080628
HEIE (ours) 1.031390 4.582199 4.352880
Table 2: Performance of image implausibility explanations on our Expl-AIGI-Eval Dataset.
All Data $GT=0$ $GT>0$
MSE ↓ MSE ↓ KLD↓ CC↑ SIM↑ AUC-Judd↑
InternViT[8] 0.07318 0.07248 3.515 0.019 0.091 0.524
EVA-CLIP[32] 0.00924 0.00207 3.226 0.582 0.095 0.607
CLIP[29] 0.00916 0.00920 1.953 0.244 0.154 0.636
HEIE (ours) 0.00510 0.00076 1.629 0.684 0.423 0.938
Table 3: Results on AbHuman. All models are finetuned.
Train: RichHF-18K All Data $GT=0$ $GT>0$ Train: Abhuman All Data $GT=0$ $GT>0$
Test: AbHuman [10] MSE (All Data)↓ MSE (GT=0)↓ KLD↓ CC↑ Test: RichHF-18K MSE (All Data)↓ MSE (GT=0)↓ KLD↓ CC↑
InternViT-300MB [8] 0.12610 0.12384 3.263 -0.046 InternViT-300MB 0.25362 0.25170 10.547 -0.017
CLIP [29] encoder 0.01181 0.00492 2.730 0.212 CLIP encoder 0.01510 0.00384 3.448 0.015
EVA-CLIP[32] 0.01170 0.00562 3.225 0.482 EVA-CLIP 0.01432 0.00148 2.835 0.414
HEIE (ours) 0.00858 0.00298 1.796 0.506 HEIE (ours) 0.01091 0.00021 1.996 0.433
Table 4: Performance of zero-shot domain generalization on the RichHF-18K and AbHuman dataset.
4 Experiments
4.1 Experimental Setup
4.1.1 Implementation Details
Our model is based on the InternVL-8B [8]. Our Expl-AIGI-Eval dataset is constructed by expanding upon two existing datasets, RichHF-18K and AbHuman, by adding explanatory and analytical annotations. We follow the original train-validation split. We use Deepspeed to finetune the MLLM, with a learning rate of $3\times 10^{-4}$, a warmup ratio of 0.03, and a batch size of 16.
4.1.2 Evaluation Metrics
We follow the method in [23] for evaluation. We report the MSE for all samples, including images with no problematic region. For images with implausibility regions, we provide evaluation metrics for the saliency heatmaps, such as KLD, AUC-Judd, SIM, and CC, as referenced in [5]. For the score prediction tasks, we report the Pearson linear correlation coefficient (PLCC) and Spearman rank correlation coefficient (SRCC) [5]. Additionally, since the RAHF model [23] is not open source and its training is complex, we can only provide the results reported in the RichHF paper and cannot test the actual performance of the RAHF model.
4.2 Comparison with State-of-the-art
We evaluate our method against the RichHF-18K benchmark [23], as presented in Tab. 1. Conventional vision models typically lack physical common sense and struggle with high-level semantic understanding, complicating their ability to perceive and recognize content that goes against human intuition. In contrast, our method excels in accurately predicting heatmaps. Given the strong correlation between heatmap accuracy and image plausibility scores, our approach provides quantitative metrics that align with human evaluations. In comparison, our model achieved SOTA performance.
As shown in Tab. 2, we evaluate the performance of image implausibility explanations on our Expl-AIGI-Eval Dataset. We employe three evaluation methods to evaluate the models’ text explanation capability: (1) Compare the perplexity (ppl) of model outputs to ground truth; (2) Employ GPT-4o to compare the content similarity between the model outputs and the ground truth. (3) Conduct human evaluations of the model-generated explanations. Our carefully designed CoT approach significantly enhances our analysis and explanation capabilities for AIGC image implausibility. Our textual analysis outperforms both other open-source MLLMs and surpasses the general-purpose SOTA models like GPT-4o and Claude-3.5-sonnet. Across various textual explanation metrics, our model achieves SOTA performance.
Tab. 3 shows the results on the AbHuman dataset [10], which focuses more on the flaws in AI-generated images of humans. Baseline models underperform due to limited common sense and reasoning skills. In contrast, our model achieves SOTA results. Additionally, the RAHF model’s performance remains uncertain due to undisclosed results and its complex, closed-source training process.
4.3 Zero-shot Domain Generalization Results
Tab. 4 presents zero-shot cross-domain generalization experiments. We train on either the RichHF-18K or AbHuman datasets and perform zero-shot inference on the other. Cross-domain tasks [48, 13], common in real-world applications [46, 47], test a model’s fundamental world knowledge beyond domain-specific fitting [12]. As shown, smaller models suffer significant accuracy drops, losing defect assessment ability. In contrast, our model excels, leveraging MLLMs’ strong common-sense knowledge and zero-shot capabilities.
4.4 Ablation Study
Exps Global Local FW LW UncW MSE↓ KLD↓ CC↑
(a) 0.01071 1.950 0.502
(b) 0.00980 1.921 0.504
(c) 0.00954 1.874 0.511
(d) 0.00873 1.680 0.557
(e) 0.00825 1.634 0.574
Table 5: Ablation study for Adaptive Hierarchical Implausibility Mapper on RichHF-18K dataset.
Ablation Study for Adaptive Hierarchical Implausibility Mapper. As shown in Tab. 5, Exp. (a) uses only a special global map token and global image features for implausibility prediction, resulting in poor accuracy due to lost local details. Exp. (b) uses only a special local map token, preserving fine-grained image information and improving accuracy over Exp. (a). Exp. (c) combines global and local tokens with fixed weights, outperforming both singular approaches. Exp. (d) employs learnable weights, further improving results over Exp. (c). Exp. (e) uses an uncertainty-based adaptive strategy, dynamically weighting local and global heatmaps and achieving optimal performance.
Exps Token Heatmap GT Heatmap PLCC↑ SRCC↑
(a) 0.216 0.197
(b) 0.653 0.640
(c) 0.597 0.575
(d) 0.697 0.683
(e) 0.743 0.710
Table 6: Ablation for Verisimilitude Scorer on RichHF-18K dataset.
Ablation for Verisimilitude Scorer. Tab. 6 presents the results: Exp. (a) outputs scores in plain text via LLM, showing low accuracy due to LLM’s numerical insensitivity. Exp. (b) uses the LLM’s special score token [SCORE] for score prediction, while Exp. (c) uses only heatmap data. Both improve significantly over plain text. Exp. (d), our Verisimilitude Scorer, combines heatmap and [SCORE] token features with a calibration function for superior accuracy, validating its effectiveness. Exp. (e) further validates using ground truth heatmaps and [SCORE] token, achieving excellent score predictions, demonstrating the importance of accurate heatmaps.
Heatmap Score
MSE↓ KLD↓ CC↑ PLCC↑ SRCC↑
w/o CoT Text 0.00913 1.689 0.553 0.669 0.658
w/ CoT Text 0.00825 1.634 0.574 0.697 0.683
w/ GT CoT Text 0.00792 1.601 0.580 0.701 0.685
Table 7: Ablation Study for CoT-Driven Explainable System.
Ablation Study for CoT-Driven Explainable System. As shown in Tab. 7, we evaluate our CoT Instruction. The first experiment without CoT Text (w/o CoT Text) reports results lacking our CoT Instruction, while the second (w/ CoT Text) shows outcomes with it. Comparing these, CoT Prompt significantly enhances heatmap and score predictions, demonstrating its effectiveness in guiding logical reasoning in LLMs. The third experiment (w/ GT CoT Text) uses ground truth CoT text labels, including image descriptions and problematic region identification, as instruction. These further boost prediction accuracy, affirming CoT’s value in enhancing LLM [45, 43] reasoning.
4.5 Visualization Analysis
Evaluation Results of HEIE: As shown in Fig. 4, by using CoT, HEIE implements comprehensive and explainable AIGC image implausibility evaluation. Our model does not produce false positive predictions for defect-free images.
Comparison with Baseline: As Fig. 5 illustrates, predicting image implausibility heatmaps is challenging, and baselines perform poorly. Our model, however, shows excellent predictive performance. As well, we cannot obtain RAHF’s prediction results due to the unavailability of RAHF’s open model.
Results of Hierarchical Implausibility Mapper: As shown in Fig. 6, global heatmaps tend to predict obvious, coarse-grained regions, while local heatmaps focus on subtler, finer details. Their adaptive combination yields the final heatmap.
5 Conclusion
In summary, we introduce the CoT-Driven Explainable Trinity Evaluator, which guides LLM to generate the heatmap, score, and explanation for image implausibility. We propose an Adaptive Hierarchical Implausibility Mapper capable of enhancing the prediction of subtle implausibilities. Additionally, we present the Expl-AIGI-Eval dataset, which offers high-quality, explainable implausibility annotations. Our work facilitates users in understanding the image implausibility and improving the images. In the future, we aim to improve the reasoning efficiency and performance of the MLLM-based framework.

### v2

@@
 Specialized models aim to predict defect region heatmaps but face two primary challenges: (1) lack of explainability, failing to provide reasons and analyses for subtle defects, and (2) inability to leverage common sense and logical reasoning, leading to poor generalization.
-Multimodal large language models (MLLMs) promise better comprehension and reasoning but face their own challenges: (1) difficulty in fine-grained defect localization due to the limitations in capturing tiny details; and (2) constraints in providing pixel-wise outputs necessary for precise heatmap generation.
+Multimodal large language models (MLLMs) promise better comprehension and reasoning but face their own challenges: (1) difficulty in fine-grained defect localization due to the limitations in capturing tiny details, and (2) constraints in providing pixel-wise outputs necessary for precise heatmap generation.
 To address these challenges, we propose HEIE: a novel MLLM-Based Hierarchical Explainable Image Implausibility Evaluator.
@@
 Our method demonstrates state-of-the-art performance through extensive experiments.
+Our project is at https://yfthu.github.io/HEIE/.
 Figure 1: (a) Specialized models lack the ability to explain and analyze subtle implausibility regions, hindering understanding for general users.
@@
 However, they often face quality issues such as artifacts, unnatural textures, and problematic regions.
-Accordingly, many studies focus on the evaluation of AIGC images [44, 42, 19, 26].
+Accordingly, many studies focus on the evaluation of AIGC images [51, 50, 22].
 Current evaluations primarily yield scalar scores, simple 0-1 values, failing to identify specific defect regions, which hinders targeted improvements in generative models.
@@
 However, specialized models like RAHF have several limitations in the implausibility evaluation task:
-  1. 1.
-    Lack of explainability.
+(1) Lack of explainability.
 As shown in Fig.
@@
 Specialized small models fail to provide detailed explanations for these subtle defects, making it difficult for general users to understand and optimize the AIGC models.
-  2. 2.
-    Lack of general knowledge.
+(2) Lack of general knowledge.
 These small specialized models are trained on limited datasets and lack comprehensive world knowledge, common sense, and logical reasoning skills.
@@
 However, directly applying vanilla MLLM to image implausibility evaluation remains challenging:
-  1. 1.
-    Fine-Grained Defect Localization: While MLLMs excel in capturing global image structure, they struggle with pinpointing fine-grained defects.
+(1) Fine-Grained Defect Localization: While MLLMs excel in capturing global image structure, they struggle with pinpointing fine-grained defects.
 Tiny defect areas like the corners of eyes or eyebrows (Fig.
 1 (b)) are often neglected by strong MLLMs such as GPT-4o.
-  2. 2.
-    Pixel-Wise Output Limitations: General-purpose MLLMs handle various modalities (text, images, videos, etc.) but usually output text only, lacking the capability to produce pixel-level defect heatmaps (Fig.
+(2) Pixel-Wise Output Limitations: General-purpose MLLMs handle various modalities (text, images, videos, etc.) but usually output text only, lacking the capability to produce pixel-level defect heatmaps (Fig.
 1 (b)).
@@
   4. 4.
-    We construct a high-quality explainable AIGC image implausibility evaluation dataset, which will be publicly available.
-Our method achieves state-of-the-art performance on various datasets and tasks. Extensive comparisons and ablation studies validate the effectiveness of our approach.
+    We construct and publish a high-quality explainable AIGC image implausibility evaluation dataset.
+Our method achieves state-of-the-art performance on various datasets and tasks.
+Extensive comparisons and ablation studies validate the effectiveness of our approach.
 Figure 2: The Implausibility Mapper processes a dynamic number of special [MAP] tokens and the image features to enhance detailed implausibility localization.
@@
 Large-scale visual understanding models are now commonly used for reliable evaluation due to their strong semantic capabilities [30, 49, 56].
-Prior research has shown that MLLMs are proficient in both low-level visual evaluations [37, 39] and high-level semantic comprehension [52, 49].
+Prior research has shown that MLLMs are proficient in both low-level visual evaluations [45, 47] and high-level semantic comprehension [58, 55, 36, 35, 34].
 Some studies  [46, 39, 20, 48] activate the visual evaluation capabilities of multimodal MLLMs through data fine-tuning.
@@
 HumanRefiner targets defects in AIGC-generated human images, while RichHF encompasses defects in various objects.
-However, these approaches rely on proprietary small models to detect defect regions.
+However, these approaches rely on specialized small models to detect defect regions.
 These models are limited in common sense and knowledge, resulting in insufficient generalization capabilities in complex scenarios.
-Additionally, proprietary small models lack interpretability and cannot provide detailed reasons or explanations  [2, 9] for the detected issues.
+Additionally, specialized small models lack interpretability and cannot provide detailed reasons or explanations  [2, 12] for the detected issues.
 3 Method
@@
 We introduced a special map token [MAP] to extract implausibility region information from LLM.
-And we use image features $F$ from the image encoder and [MAP] token features $T$ from the LLM for a two-way cross-attention calculation [18, 1, 35], as shown in Eq.
+In LLM, the [MAP] token generates many feature layers, and its final linear layer feature is used to predict the next token for LLM.
+We select this final LLM layer’s hidden state of the [MAP] token as the features $T$.
+Then we use image features $F$ from the image encoder and [MAP] token features $T$ from the LLM for a two-way cross-attention calculation [21, 1, 40], as shown in Eq.
 LABEL:2waytrans.
-To fuse the [MAP] token embeddings with image features, each input alternates as the query and value during cross-attention operations.
-Specifically, The special map token embeddings $T$ and image features $F$ server as $E_{\text{text}}$ and $E_{\text{img}}$ respectively in the following two-way cross-attention:
-$C_{\text{text\textrightarrow img}}=\text{CrossAtten}(\text{SelfAtten}(E_{\text{text}}),E_{\text{img}},E_{\text{img}})$ (1)
-$E_{\text{text}^{{}^{\prime}}}=\text{mlp}(\text{norm}(C_{\text{text\textrightarrow img}}))$
-$C_{\text{img\textrightarrow text}}=\text{CrossAtten}(E_{\text{img}},E_{\text{text}^{{}^{\prime}}},E_{\text{text}^{{}^{\prime}}})$
-$O_{2way\_CA}=\text{norm}(C_{\text{img\textrightarrow text}})$
-By stacking the above two-way cross-attention twice, we get the predicted heatmap $H$.
+$C_{\text{T\textrightarrow F}}=\text{CrossAtten}(\text{SelfAtten}(T),F,F)$ (1)
+$T^{\prime}=\text{mlp}(\text{norm}(C_{\text{T\textrightarrow F}}))$
+$C_{\text{F\textrightarrow T}}=\text{CrossAtten}(F,T^{\prime},T^{\prime})$
+$O_{2way\_CA}=\text{norm}(C_{\text{F\textrightarrow T}})$
+To fuse the [MAP] token features $T$ with image features $F$, each input alternates as the query and value during cross-attention operations.
+By stacking the above two-way cross-attention Eq.
+(LABEL:2waytrans) twice, we get the predicted heatmap $H$.
 Data imbalance poses a challenge in heatmap prediction, with mean squared error (MSE) loss often leading to conservative and low-value predictions.
@@
 The LLM is guided through a Chain of Thought (CoT) prompting process in five steps:
-1. Image Description: The LLM describes the image, capturing key elements in natural language.
-2. Problematic Regions Identification: Based on the description and the understanding of the image, the LLM identifies potential problematic areas for further analysis.
-3. Special map token: Based on the preceding Problematic Regions Identification, the LLM injects defect region information into the special map token for use by the implausibility mapper.
-4. Analysis of Issues: Based on the preceding implausibility localization, the LLM provides detailed text explanations for each region, including type, causes, and so on.
-5. Special Score Token: Based on the comprehensive understanding, the LLM incorporates overall information into a special score token for use by our Verisimilitude Scorer.
+(1) Image Description: The LLM describes the image, capturing key elements in natural language.
+(2) Problematic Regions Identification: Based on the description and the understanding of the image, the LLM identifies potential problematic areas for further analysis.
+(3) Special map token: Based on the preceding Problematic Regions Identification, the LLM injects defect region information into the special map token for use by the implausibility mapper.
+(4) Analysis of Issues: Based on the preceding implausibility localization, the LLM provides detailed text explanations for each region, including type, causes, and so on.
+(5) Special Score Token: Based on the comprehensive understanding, the LLM incorporates overall information into a special score token for use by our Verisimilitude Scorer.
 This CoT-driven process breaks down the complex evaluation task into five progressively challenging and interrelated tasks, fully leveraging the LLM’s potential.
@@
 We provide detailed analyses and explanations for each implausibility region.
-The annotation process is conducted in three stages, shown in Fig.
+The annotation process is conducted in the following three stages, shown in Fig.
 3.
-1.
-Visual Prompting: Research [6] indicates that LLMs effectively adapt to visual prompts.
-Consequently, we annotate defect regions directly on images, ensuring visibility while preserving the image’s integrity and facilitating subsequent use by Claude-3.5-sonnet.
-2.
-LLM Free-Form Outputting: Using Claude-3.5-sonnet, we generate detailed analyses from visual prompts.
-Prior studies [33] reveal that fixed-format enforcement degrades performance, so we allow free-form outputs for defect causes and analyses.
-3. In-Context Learning-Based Formatting: To ensure consistency, we employ in-context learning to guide the GPT-4o in refining and formatting information.
+(1) Visual Prompting: We follow [7] to add visual prompts.
+Specifically, the original dataset provides masks of the defective regions for each image.
+We draw a red contour for each defective region, clearly highlighting these areas, as shown in Stage 1 in Fig.
+3.
+(2) LLM Free-Form Outputting: Using Claude-3.5-sonnet [3], we generate detailed analyses of the defect regions.
+Previous study [38] indicates that enforcing fixed-format outputs impairs the performance of LLM, so we allow the LLM to produce free-form outputs for analyzing defect causes.
+(3) In-Context Learning-Based Formatting: As shown in Stage-3 (Fig.3), we employ GPT-4o and in-context learning to format the defect analysis texts obtained in Stage 2.
 This hierarchical approach ensures comprehensive annotations.
@@
 3.3.2 and aiding CoT training.
-In summary, our Expl-AIGI-Eval dataset comprises 42,760 AIGC images divided into two parts: RichHF-18K [23] and AbHuman [10].
-We create 20 diverse query expressions per image for LLM instructions.
-We manually review 20% of the generated content for quality assurance.
 Sensitive images related to pornography, privacy, animal protection, and political topics are filtered out to avoid ethical issues.
@@
 For the score prediction tasks, we report the Pearson linear correlation coefficient (PLCC) and Spearman rank correlation coefficient (SRCC) [6].
-Additionally, since the RAHF model [23] is not open source and its training is complex, we can only provide the results reported in the RichHF paper and cannot test the actual performance of the RAHF model.
+Additionally, since the RAHF model [26] is not open source, we provide its results reported in its paper.
 4.2 Comparison with State-of-the-art
-We evaluate our method against the RichHF-18K benchmark [23], as presented in Tab.
+We evaluate our method on the RichHF-18K benchmark [26], as presented in Tab.
 1.
@@
 2, we evaluate the performance of image implausibility explanations on our Expl-AIGI-Eval Dataset.
-We employe three evaluation methods to evaluate the models’ text explanation capability: (1) Compare the perplexity (ppl) of model outputs to ground truth; (2) Employ GPT-4o to compare the content similarity between the model outputs and the ground truth.
+We employ three evaluation methods to evaluate the models’ text explanation capability: (1) Compare the perplexity (ppl) of model outputs to ground truth; (2) Employ GPT-4o to compare the content similarity between the model outputs and the ground truth.
 (3) Conduct human evaluations of the model-generated explanations.
@@
 In contrast, our model achieves SOTA results.
-Additionally, the RAHF model’s performance remains uncertain due to undisclosed results and its complex, closed-source training process.
 4.3 Zero-shot Domain Generalization Results
@@
 We train on either the RichHF-18K or AbHuman datasets and perform zero-shot inference on the other.
-Cross-domain tasks [48, 13], common in real-world applications [46, 47], test a model’s fundamental world knowledge beyond domain-specific fitting [12].
+Cross-domain tasks [54, 16], common in real-world applications [42, 41, 11, 10, 53, 52], test a model’s fundamental world knowledge beyond domain-specific fitting [15, 43].
 As shown, smaller models suffer significant accuracy drops, losing defect assessment ability.
@@
 The third experiment (w/ GT CoT Text) uses ground truth CoT text labels, including image descriptions and problematic region identification, as instruction.
-These further boost prediction accuracy, affirming CoT’s value in enhancing LLM [45, 43] reasoning.
+These further boost prediction accuracy, affirming CoT’s value in enhancing LLM reasoning.
 4.5 Visualization Analysis